OpenStudy (anonymous):

I need homework help. I will give a medal to someone kind enough to help me.

OpenStudy (anonymous):

what do you need help on?

OpenStudy (anonymous):

The graph of a linear function is shown on the grid below. The graph has a diagnal line with two points. 0,-6 and 7,8

OpenStudy (anonymous):

I need help figuring out what function represents it

OpenStudy (anonymous):

A y =1/2 x − 6 B y =1/2 x + 3 C y = 2x + 3 D y = 2x − 6

OpenStudy (anonymous):

those are my answer choses